Using AVR write a subroutine that can create a time delay that is a multiple of 1ms. The multiple is passed in register r16, assume the CPU clock in 16MHz. (please include full working code) The processor used is the ATMEGA328P
Answer #1

While planning defer programs in (XXXX microcontroller), ascertaining the underlying worth that must be stacked inot TH and TL registers frames a critical thing. Give us a chance to perceive how it is finished.

Expect the processor is timed by a 16MHZ precious stone.

That implies, the clock info will be 16MHZ/16 = 1MHz

That implies, the time taken for the clock to make one augmentation = 1/1MHz = 1uS

For a period postponement of "X" uS the clock needs to make "X" increases.

2^16 = 65536 is the maximim number of checks feasible for a 16 bit clock.

Give TH a chance to be the esteem that must be stacked to TH registed and TL be the esteem that must be stacked to TL enlist.

At that point, THTL = Hexadecimal likeness (65536-X) where (65536-X) is considered in decimal.

Illustration.

Give the required defer a chance to be 1000uS (ie; 1mS).

That implies X = 10001

65536 – X = 65536 – 1000 = 64536.

64536 is considered in decimal and changing over it t0 hexadecimal gives FC18

That implies THTL = FC18

Along these lines TH=FC and TL=18

Program for producing 1mS defer utilizing XXXX microcontroller clock.

The program appeared beneath can be utilized for producing 1mS deferral and it is composed as a subroutine with the goal that you can call it anyplace in the program. Likewise you can place this in a circle for making longer time delays (products of 1mS). Here Timer 0 of XXXX microcontroller is utilized and it is working in MODE1 (16 bit clock).

Deferral: MOV TMOD,#00000001B/Sets Timer 0 to MODE1 (16 bit clock). Clock 1 isn't utilized

MOV TH0,#0FCH/Loads TH0 enroll with FCH

MOV TL0,#018H/LOads TL0 enroll with 18H

SETB TR0/Starts the Timer 0

HERE: JNB TF0,HERE/Loops here until the point when TF0 is set (ie;until move over)

CLR TR0/Stops Timer 0

CLR TF0/Clears TF0 hail

RET

The above defer routine can be circled twice so as to get a 2mS deferral and it is appeared in the program beneath.

Primary: MOV r16,#2D/number of ms you need delay

Circle: ACALL DELAY

DJNZ r16,LOOP

SJMP MAIN

Postponement: MOV TMOD,#00000001B

MOV TH0,#0FCH

MOV TL0,#018H

SETB TR0

HERE: JNB TF0,HERE

CLR TR0

CLR TF0

RET